OpenStudy (anonymous):

A total of 282 tickets were sold for the school play. They were either adult tickets or student tickets. The number of student tickets sold was two times the number of adult tickets sold. How many adult tickets were sold?

OpenStudy (anonymous):

I'm guessing is 282 divided by 2 first?

OpenStudy (beccaboo333):

I'm confused on this one. @iambatman will help you.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

...

OpenStudy (anonymous):

???

OpenStudy (anonymous):

x = # of adult tickets y = # of student tickets sold x+y = 282 y = 2x x +(2x) = 282 3x = 282 x = 94 x + y = 282 92 + y = 282 y = 282 - 92 y = 190 Therefore 92 adult tickets were sold, and 190 student tickets were sold.

OpenStudy (anonymous):

it says it's wrong

OpenStudy (anonymous):

Btw, you should be more respectful next time, rather than putting... and question marks. Have some patience.

OpenStudy (anonymous):

Sorry I have like 5 minutes to do this ( I have online school)

OpenStudy (anonymous):

It's not wrong

OpenStudy (anonymous):

I put in 92

OpenStudy (anonymous):

Ok, so what else might it be?

OpenStudy (anonymous):

No, there's nothing wrong with what I wrote.

OpenStudy (anonymous):

oh the answer is 94 not 92

OpenStudy (anonymous):

I put 92 in by accident for second part it should be 94

OpenStudy (anonymous):

Yeah thats right thanks

OpenStudy (anonymous):

But, yes your final answer is 94 :)

OpenStudy (anonymous):

Do you have time for like 1 or 2 more?

OpenStudy (anonymous):

Well, I have to go finish my assignment right now. Maybe later tomorrow if you're on.

OpenStudy (anonymous):

Oh ok, probably not, but thanks anyway man!

OpenStudy (anonymous):

Np :)